1.Let f(x) = (x - 8)2. Find x so that f(x) = 100
2.Write a quadratic equation having the given numbers as solutions.
3, -8

Answer by stanbon(75887)

1.Let f(x) = (x - 8)^2. Find x so that f(x) = 100
---
100 = (x-8)^2
Rearrange:
(x-8)^2-100 = 0
Factor:
(x-8-10)(x-8+10) = 0
(x-18)(x+2) = 0
x = 18 or x = -2
===================
2.Write a quadratic equation having the given numbers as solutions.
3, -8
---
f(x) = (x-3)(x+8)
f(x) = x^2 + 5x - 24
==========================
Cheers,
Stan H.